A PRIMER ON DUNHAM’S APPROACH TO ANALYSIS OF SPECTRA OF FREE DIATOMIC MOLECULES

\underline{J. F. OGILVIE}, Escuela de Quimica, Universidad de Costa Rica, San Pedro, San Jose, Costa Rica, and Centre for Experimental and Constructive Mathematics, Department of Mathematics, Simon Fraser University, 8888 University Drive, Burnaby, British Columbia V5A 1S6 Canada.